OpenStudy (anonymous):

To open these doors, you must match the number and type of solutions for the following two fuctions in standard form. f(x) = x2 + 6x – 16 g(x) = x2 +6x + 1

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@hazem-wardat

OpenStudy (jdoe0001):

have you covered factoring of quadratics yet?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

well kinda but i dont get this stuff

OpenStudy (jdoe0001):

well, the 1st one is very straightforward the 2nd one, no so much I'd assume you haven't covered the quadratic formula yet?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

f(x) = x2 + 6x – 16 x1= 2 x2=-3 for g(x) = x2 +6x + 1 x1= \[-3+2\sqrt{2}\] x2=\[-3-2\sqrt{2}\]

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

hhhhhh it's my game

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=eF6zYNzlZKQ <--- factoring quadratics